A square poster has sides measuring 2 feet less than the sides of a square sign if the difference between their areas is 12 square feet find the lengths of the sides of the poster and sign
Answer by ankor@dixie-net.com(22740)

A square poster has sides measuring 2 feet less than the sides of a square sign
if the difference between their areas is 12 square feet find the lengths of the sides of the poster and sign
:
let x = the side of the poster
then
(x+2) = the side of the sign
:
Sign area - poster area = 12
(x+2)^2 - x^2 = 12
FOIL(x+2)(x+2)
x^2 + 2x + 2x + 4 - x^2 = 12
combine like terms
4x + 4 = 12
4x = 12 - 4
4x = 8
x = 8/4
x = 2 ft is the side of the poster, then 4 ft is the side of the sign
:
:
Check this
4^2 - 2^2 =
16 - 4 = 12